OpenStudy (anonymous):

I really need help with this, I will medal :D A line has a slope of 3 and contains the point (−1, −8). Which equations represent the line? A. −3x + y = −5 B. 3x + y = −5 C. (y – 8) = 3(x − 1) D. (y + 8) = 3(x + 1)

OpenStudy (anonymous):

I think its A and D

OpenStudy (welshfella):

the point slope form of a line is y - y1 = m(x - x1) where m = slope and (x1,y1) is a point on the line plug in the given values to get your equation

OpenStudy (anonymous):

so its A and D....?

OpenStudy (welshfella):

m = 3 , x1 = -1 and y1 = -8

OpenStudy (welshfella):

plugging them in gives option D Yes

OpenStudy (anonymous):

ok, so what about option a?

OpenStudy (welshfella):

now you need to rearrange equation D and see if you get one of the other options

OpenStudy (welshfella):

y + 8 = 3(x +1) y + 8 = 3x + 3 can you continue?

OpenStudy (anonymous):

I still get option a...

OpenStudy (anonymous):

:( I think its A and D...

OpenStudy (welshfella):

yes you are correct

OpenStudy (anonymous):

thanks....can you help with another? or are you busy...? XD I will still medal you though

OpenStudy (welshfella):

y + 8 = 3x + 3 subtract 3x from both sides -3x + y + 8 = 3 subtrcat 8 from both sides -3x + y = -5

OpenStudy (anonymous):

Which graph is the line y=-(1/5x)-1 ?

OpenStudy (welshfella):

no i have about 15 minutes but please re post

OpenStudy (welshfella):

oh ok carry on

OpenStudy (anonymous):

ok, thanks for the help, I will tag you